# Singularity of Lévy walks in the lifted Pomeau-Manneville map

### v1

@@ IV.1 Results
Figure 5: The GDC $K(z)$ (2) generated from TGK expansions (II.1) of the superdiffusive PM map (3), using increasingly many terms $t$ of the expansion and the analytic estimate for $\beta$ obtained from (11) (rough curves, $t$ increasing from light to dark). Also shown is the LW approximation taken from (11) and (2) (smooth curve, red online). The different dynamical regimes shown in (11) are annotated on the plotting area. All plots are generated from an ensemble of $10^{4}$ simulated trajectories and consist of $151$ points.
@@ IV.2 Understanding the discrepancy between deterministic dynamics and CTRW theory
Figure 6: $K(z)$ generated from TGK expansions (II.1) of the superdiffusive PM map, using increasingly many orders of correlation terms $\langle v_{k}v_{k+j}\rangle$ (lower rough curves, $m$ increasing from dark to light). Also shown are the LW approximation (11) (uppermost smooth curve, red online) and the full TGK expansion (uppermost rough curve, blue online). All TGK plots are taken using $200$ terms of a TGK expansion from an ensemble of $10^{4}$ simulated trajectories and consist of $101$ points.
@@
Figure 7: Histograms of the time spent in the left-hand branch (adjacent to the marginal point) of the PM map, calculated from pre-iterates of the map, for varying values of $z$. In red for comparison is shown the corresponding density $w(t)$ as prescribed by (9). Backwards iterates from the map are generated using numerical root finding using Python’s decimal package, to 40 decimal places.
@@ IV.3 Effect of aging on the generalised diffusion coefficient
Figure 8: Top: TGK approximations of $K(z)$ for the superdiffusive PM map, starting from increasingly aged ensembles of particles initialised from a uniform density (aging duration increasing from light to dark), as well as starting from a normalisable invariant density (where it exists; darkest uppermost curve, blue online). Also shown are the results for the non-aged LW discussed above (smooth curve, red online) and the aged LW discussed in Albers and Radons (2018, 2022) (dashed line, top, red online). All TGK plots are taken using $200$ terms of a TGK expansion from an ensemble of $10^{4}$ simulated trajectories and consist of $151$ points ($50$ points in the case of the invariant ensemble). Bottom left: an identical plot to that shown in Fig. 5, but with points initialised from an ensemble aged by $100$ timesteps. Bottom right: an identical plot to that shown above, for the subdiffusive extension of the map studied in Korabel et al. (2007), mirroring Fig. 7 in that paper, with $a=5$ fixed (aging duration increasing from dark to light, invariant curve shown at the bottom, in teal online).
@@
Finally, we explore the consequences of aging on numerical estimates of $K(z)$. In the present setting, aging denotes the dependence of the MSD, and hence of $K(z)$, on the duration, or aging time, between the initialisation of the system dynamics and the start of its measurement Bouchaud (1992); Barkai (2003); Barkai and Margolin (2004); Akimoto and Barkai (2013); Metzler et al. (2014b). Up to this point all estimates have been taken commencing from a uniform density of $\{x_{0}\}$. However, many systems observed in nature typically evolve for a long time before human measurement of them can commence, and therefore we also consider the case where we allow a degree of relaxation towards the invariant density to take place before beginning measurements. We denote the number of timesteps during which the system is allowed to evolve before measurement begins as the ‘aging time’ $t_{a}$. In such cases it is known that the dynamics of systems exhibiting aging are characteristically dependent on the ratio of measurement time to aging time, $t/t_{a}$ Akimoto and Barkai (2013).
@@
In Fig. 8 we apply varying degrees of aging to an ensemble of trajectories before calculating $K(z)$ via a numerical TGK expansion. We also show the same calculations starting from the (approximate) invariant density where it exists for $z<2$; inspired by Thaler (1983); Zweimüller (1998); Thaler (2000); Korabel and Barkai (2010); Akimoto and Barkai (2013), which show the invariant density of the reduced map (1) to be asymptotic to $\tilde{\rho}(x)\propto x^{1-z}$, we initialise our points from
$\tilde{\rho}(x)\propto x^{1-z}+(1-x)^{1-z}.$ (15)
@@
For $z\geq 2$ the only normalisable invariant density is the delta function on $\{0,1\}$, which implies almost all particles undergo purely ballistic motion in a single direction. Also shown is the corresponding analytic result for a LW in the limit of large aging relative to measurement time, $t/t_{a}\ll 1$, as a special case of recently reported results in Albers and Radons (2022). It is known that the MSD of aged walks corresponds exactly to the ensemble-averaged time-averaged MSD, with the averaging time equated to the aging time $t_{a}$ Bothe et al. (2019). In Albers and Radons (2022) this was calculated for a generalised LW with non-constant velocity, of which our stochastic model is merely a special case.
@@
That there is no effect of aging on $K(z)$ in regime (I) for $z<\frac{3}{2}$ is to be expected, since in this regime, any initial density converges to the invariant one in exponential time. In contrast, it is known that for $z\geq 2$ the PM map is only weakly ergodic Akimoto and Barkai (2013); Metzler et al. (2014a), and therefore undergoes dynamical aging Akimoto and Barkai (2013); Korabel et al. (2007); Zweimüller (1998); Thaler (1983). Interestingly though, some aging is observed on $K(z)$ even in the intermediate superdiffusive regime (II), $\frac{3}{2}<z<2$, in which a normalisable invariant density still exists. This reflects what was described for LWs as ‘ultraweak’ ergodicity breaking in that parameter regime, which is characterised by a deviation in the GDC, but not the exponent, of the time dependence of the MSD Metzler et al. (2014c); Froemberg and Barkai (2013a, b).
@@
We conclude this section by presenting a general principle: that, for measurements taken over finite time, we believe it is (perhaps counterintuitively) preferable to average one’s ensembles over a uniform initial density than from one closer to the more ‘natural’ invariant density. This is especially evident in those regimes (II) and (III) for which $z>\frac{3}{2}$, in which the map’s ergodicity is broken. We see that, this way, results from a uniform initial density in practice better reproduce the curve seen in the long measurement time limit, without simulations becoming dominated by effects introduced by aging; and for short sampling times, measuring from an aged or even invariant density can produce perverse results which are not representative of the long-measurement-time GDC of the process. Our exposition here supplies some numerical evidence to back this principle.
@@ V Conclusion
We also investigated the effect of aging on our numerical findings. We find, in accordance with recently reported results, that in the limit of increased aging time $t_{a}/t$, the suppression of $K(z)$ at the transition from superdiffusion to ballistic motion is eliminated, for finite measurement times, while the divergence at the other, from normal to superdiffusion, is preserved. This elimination is also reproduced in the case of the subdiffusive PM map. We identify this as a problem of conflicting limits as both aging time and measurement time go to infinity, which requires one to take caution when conducting numerical or even physical experiments. We state a general principle that uniform initial densities should be preferred in simulations to more ‘natural’ invariant ones if the ergodicity of the dynamics is not ensured, a claim which is supported numerically.
@@
Appendix
V.1 CTRW calculations
@@ V.2 Bifurcation calculations
[…]

### v2

@@ IV.1 Results
Figure 5: The GDC $K(z)$ (2) generated from TGK expansions (II.1) of the superdiffusive PM map (3), fixing $R=1$, using increasingly many terms $t$ of the expansion and the analytic estimate for $\beta$ obtained from (11) (rough curves, $t$ increasing from light to dark). Also shown is the LW approximation taken from (11) and (2) (smooth curve, red online). The different dynamical regimes shown in (11) are annotated on the plotting area. All plots are generated from an ensemble of $10^{4}$ simulated trajectories and consist of $151$ points.
@@ IV.2 Understanding the discrepancy between deterministic dynamics and CTRW theory
Figure 6: $K(z)$ generated from TGK expansions (II.1) of the superdiffusive PM map, fixing $R=1$, using increasingly many orders of correlation terms $\langle v_{k}v_{k+j}\rangle$ (lower rough curves, $m$ increasing from dark to light). Also shown are the LW approximation (11) (uppermost smooth curve, red online) and the full TGK expansion (uppermost rough curve, blue online). All TGK plots are taken using $200$ terms of a TGK expansion from an ensemble of $10^{4}$ simulated trajectories and consist of $101$ points.
@@
Figure 7: Histograms of the time spent in the left-hand branch (adjacent to the marginal point) of the PM map, calculated from pre-iterates of the map, for varying values of $z$, fixing $R=1$. In red for comparison is shown the corresponding density $w(t)$ as prescribed by (9). Backwards iterates from the map are generated using numerical root finding using Python’s decimal package, to 40 decimal places.
@@ IV.3 Effect of aging on the generalised diffusion coefficient
Figure 8: Top: TGK approximations of $K(z)$ for the superdiffusive PM map, fixing $R=1$, starting from increasingly aged ensembles of particles initialised from a uniform density (aging duration increasing from light to dark), as well as starting from a normalisable invariant density (where it exists; darkest uppermost curve, blue online). Also shown are the results for the non-aged LW discussed above (smooth curve, red online) and the aged LW discussed in Albers and Radons (2018, 2022) (dashed line, top, red online). All TGK plots are taken using $200$ terms of a TGK expansion from an ensemble of $10^{4}$ simulated trajectories and consist of $151$ points ($50$ points in the case of the invariant ensemble). Bottom left: an identical plot to that shown in Fig. 5, but with points initialised from an ensemble aged by $100$ timesteps. Bottom right: an identical plot to that shown above, for the subdiffusive extension of the map studied in Korabel et al. (2007), mirroring Fig. 7 in that paper, with $a=5$ fixed (aging duration increasing from dark to light, invariant curve shown at the bottom, in teal online).
@@
Finally, we explore the consequences of aging on numerical estimates of $K(z)$. In the present setting, aging denotes the dependence of the MSD, and hence of $K(z)$, on the duration, or aging time, between the initialisation of the system dynamics and the start of its measurement Bouchaud (1992); Barkai (2003); Barkai and Margolin (2004); Akimoto and Barkai (2013); Metzler et al. (2014a). Up to this point all estimates have been taken commencing from a uniform density of $\{x_{0}\}$. However, many systems observed in nature typically evolve for a long time before human measurement of them can commence, and therefore we also consider the case where we allow a degree of relaxation towards the invariant density to take place before beginning measurements. We denote the number of timesteps during which the system is allowed to evolve before measurement begins as the ‘aging time’ $t_{a}$. In such cases it is known that the dynamics of systems exhibiting aging are characteristically dependent on the ratio of measurement time to aging time, $t/t_{a}$ Akimoto and Barkai (2013).
@@
In Fig. 8 we apply varying degrees of aging to an ensemble of trajectories before calculating $K(z)$ via a numerical TGK expansion. We also show the same calculations starting from the invariant density of the reduced superdiffusive map where it exists for $z<2$, which is given in Thaler (1983); Akimoto (2012) to be
$\tilde{\rho}(x)\propto x^{1-z}+(1-x)^{1-z}$ (15)
@@
(see also discussion in Zweimüller (1998); Thaler (2000); Korabel and Barkai (2010); Akimoto and Barkai (2013), where the invariant density of the reduced map (1) is shown to be asymptotic to $\tilde{\rho}(x)\propto x^{1-z}$). For $z\geq 2$ the only normalisable invariant density is the delta function on $\{0,1\}$, which implies almost all particles undergo purely ballistic motion in a single direction. Also shown is the corresponding analytic result for a LW in the limit of large aging relative to measurement time, $t/t_{a}\ll 1$, as a special case of recently reported results in Albers and Radons (2022). It is known that the MSD of aged walks corresponds exactly to the ensemble-averaged time-averaged MSD, with the averaging time equated to the aging time $t_{a}$ Bothe et al. (2019). In Albers and Radons (2022) this was calculated for a generalised LW with non-constant velocity, of which our stochastic model is merely a special case.
@@
That there is no effect of aging on $K(z)$ in regime (I) for $z<\frac{3}{2}$ is to be expected, since in this regime, any initial density converges to the invariant one in exponential time. In contrast, it is known that for $z\geq 2$ the PM map is only weakly ergodic Akimoto and Barkai (2013); Metzler et al. (2014a), and therefore undergoes dynamical aging Akimoto and Barkai (2013); Korabel et al. (2007); Zweimüller (1998); Thaler (1983). Interestingly though, some aging is observed on $K(z)$ even in the intermediate superdiffusive regime (II), $\frac{3}{2}<z<2$, in which the map is still ergodic, and a normalisable invariant density still exists. This reflects what was described for LWs as ‘ultraweak’ ergodicity breaking in that parameter regime, which is characterised by a deviation in the GDC, but not the exponent, of the time dependence of the MSD Metzler et al. (2014a); Froemberg and Barkai (2013a, b). The effect of varying initial conditions in regime II was also discussed in Akimoto (2012) in relation to a similar superdiffusive Pomeau-Manneville-type map with bias, where a strong distributional response was identified on the MSD.
@@
We conclude this section by presenting a general principle: that, for measurements taken over finite time, we believe it is (perhaps counterintuitively) preferable to average one’s ensembles over a uniform initial density than from one closer to the more ‘natural’ invariant density. This is especially evident in regime (III), where $z>2$, for which the map is no longer ergodic, but also in regime (II) where $z>\frac{3}{2}$, where a strong dependence on initial conditions is still observed. We see that, this way, results from a uniform initial density in practice better reproduce the curve seen in the long measurement time limit, without simulations becoming dominated by effects introduced by aging; and for short sampling times, measuring from an aged or even invariant density can produce perverse results which are not representative of the long-measurement-time GDC of the process. Our exposition here supplies some numerical evidence to back this principle.
@@ V Conclusion
We also investigated the effect of aging on our numerical findings. We find, in accordance with recently reported results, that in the limit of increased aging time $t_{a}/t$, the suppression of $K(z)$ at the transition from superdiffusion to ballistic motion is eliminated, for finite measurement times, while the divergence at the other, from normal to superdiffusion, is preserved. This elimination is also reproduced in the case of the subdiffusive PM map. We identify this as a problem of conflicting limits as both aging time and measurement time go to infinity, which requires one to take caution when conducting numerical or even physical experiments. We state a general principle that uniform initial densities should be preferred in simulations to more ‘natural’ invariant ones if adequately timely mixing of the dynamics is not ensured, a claim which is supported numerically.
@@
Appendix A CTRW calculations
@@ Appendix B Bifurcation calculations
$=R-4R^{3}-64R^{7}.$
Then $x_{u}>0$ appears at $M^{2}(x_{u})=M^{-}(M^{+}(x_{u}))=x_{u}$, ie.
$(x_{u}+4Rx_{u}^{2}-R)-4R(x_{u}+4Rx_{u}^{2}-R)^{2}+R=x_{u}$
$\implies(-64R^{3})x_{u}^{4}+(-32R^{2})x_{u}^{3}+(32R^{3})x_{u}^{2}$
$\qquad\qquad\qquad+(8R^{2})x_{u}+(-4R^{3}-R)=0$
$\implies x_{u}=\frac{-1+\sqrt{1+4R^{2}}}{4R}.$
Note that $x_{u}$ is the solution of a quartic equation whose coefficients vary in $R$; hence for increasing $z$ these quantities become very difficult to calculate by hand. The critical point $R_{c}$ at which these coincide is given by the sole real positive solution to a polynomial equation of degree 7, which we solve here only numerically:
$R_{c}-8R_{c}^{3}-64R_{c}^{7}$ $=0$
$\implies R_{c}$ $\approx 0.337,$
$x_{c}$ $\approx 0.153.$
Extending these calculations to higher, or to non-integer, values of $z$ would undoubtedly lead to quite unwieldy equations, hence values of $z\in\{1,2\}$ were selected here, although the qualitative dynamics can be seen numerically to be similar.
Figure 9: Plots of sample trajectories in the Pomeau-Manneville map (mod $1$), over the region $[-\frac{1}{2},\frac{1}{2}]$, for $z=2$, $R\in\{1.15,1.25\}$. The repelling region of the phase space becomes accessible as $R$ exceeds $R_{d}\approx 1.207$, see text. The critical points ±$\pm x_{d}$ are shown in blue.
In Fig. 9, we see that the repelling region of phase space becomes accessible when the branch edge (located near $x_{n}=0$) exceeds the critical point $x_{d}$, which is a (mod $1$) repelling fixed point. In the full map, this is the point where $M(x)=x\pm 1$. Precisely, when $z=2$, this is the point
$M^{+}(x_{d})=x_{d}+4Rx_{d}^{2}-R$ $=x_{d}-1$
$\implies x_{d}$ $=\sqrt{\frac{R-1}{4R}}.$
The critical value $R_{d}$ is the value of $R$ at which $M^{-}(0)-1$ first exceeds $x_{d}$, ie.,
$M^{-}(0)-1=R_{d}-1=\sqrt{\frac{R_{d}-1}{4R_{d}}}=x_{d}$
$\implies 4R_{d}^{2}-4R_{d}+1=0$
$\implies R_{d}=\frac{1+\sqrt{2}}{2}\approx 1.207.$
Note that this is remarkably close to, but does not quite fit with, the numerically identified values for $\mathcal{N}_{1}(z)$ (the values of $R$ at which normal diffusion is restored) depicted in Fig. 2. Whether this is due to artefacts in the numerical simulations, or to something more subtle dynamically, remains an open question.
Figure 10: Zooms of Fig. 4 for $z\in\{1,2,3\}$ (left to right), for three varying levels of zooming (top to bottom), indicating the hierarchy of bubble-like structures in the neighbourhood of $R=0$, $x=0$. We conjecture this hierarchy to be infinite.